# Probabilistically Tightened Linear Relaxation-based Perturbation Analysis for Neural Network Verification

### v1

@@ 5 Empirical Evaluation
What is the general impact of PT-LiRPA in the verification process of challenging instances such as the one employed in the VNN-COMP [32, 8]?
All data are collected on a cluster running Rocky Linux 9.34 equipped with Nvidia RTX A6000 (48 GiB) and a CPU AMD Epyc 7313 (16 cores).
@@
To address the first question, we consider the original pre-trained models on MNIST dataset. Specifically, we focus on MLP models with varying depths and activation functions. For consistency and ease of comparison, we adopt the same notation as [54] and [48]: a model is denoted by the dataset name, followed by the number of layers $i$, the number of neurons per layer $j$, formatted as $i\times[j]$, and the activation function used. We begin by analyzing the mean percentage error in estimating the lower bounds of the intermediate reachable sets (using Eq. 11) for a fixed input image, using PT-LiRPA on various neural networks trained on the MNIST dataset.
@@
Name model # samples mean distance error (%) # samples mean distance error (%) # samples mean distance error (%)
1-p=0.99 1-p=0.99 1-p=0.99
MNIST 2×$\times[1024]$, ReLU 10000 12.91 100000 9.35 350000 8.78
MNIST 3×$\times[1024]$, ReLU 10000 9.63 100000 8.49 350000 7.09
MNIST $\times[1024]$, ReLU 10000 11.89 100000 9.34 350000 8.85
MNIST 2×$\times[1024]$, Tanh 10000 11.31 100000 12.22 350000 10.49
MNIST 3×$\times[1024]$, Tanh 10000 10.67 100000 11.01 350000 10.34
MNIST $\times[1024]$, Sigmoid 10000 5.03 100000 6.21 350000 4.36
@@
Table 1: Mean maximum error in estimating the lower bound of the intermediate reachable set for a fixed input image, using PT-LiRPA on various neural networks trained on the MNIST dataset with different sample sizes.
@@
The results of Tab 1 demonstrate that with high confidence (i.e., at least 99%) across all tested networks, increasing the sample size, the percentage error in the estimation decreases. Importantly, even with a limited sample size (e.g., $10000$ samples), we note that the mean error across all the networks in the lower bound estimation is $10.24\%$ from the estimated one, while increasing the sample size, we reach a maximum error of $8.32\%$. To assess the practical impact of the estimation error on the final output bounds, we consider the same perturbed input of the previous experiments and fix the sample size to $350k$.
Output Node MIP (Exact) $\alpha$-CROWN $\alpha$-CROWN w/ PT-LiRPA Soundness Tightness
1 [-17.26, -15.49] [-17.77, -15.2] [-17.44, -15.38]
2 [36.94, 39.06] [36.38, 39.45] [36.66, 39.33]
3 [-6.64, -4.39] [-7.17, -3.83] [-6.92, -4.11]
4 [3.39, 5.95] [3.12, 6.56] [3.26, 6.1]
5 [3.18, 5.61] [2.53, 5.97] [2.94, 5.74]
6 [-14.97, -12.45] [-15.37, -11.92] [-15.12, -12.26]
7 [-10.65, -8.49] [-11.08, -8.12] [-10.85, -8.29]
8 [12.11, 14.34] [11.72, 14.82] [11.98, 14.59]
9 [-4.97, -2.91] [-5.31, -2.5] [-5.08, -2.77]
10 [-17.94, -15.24] [-18.31, -14.72] [-18.07, -15.03]
Table 2: Comparison of output bounds on the MNIST_2x[1024]_ReLU network using $\alpha$-CROWN [52],$\alpha$-CROWN [52] with PT-LiRPA, and exact MIP verification [42].
We then compare the final output bounds obtained by an exact MIP-based verification [42], the state-of-the-art $\alpha$-CROWN [52] method, and $\alpha$-CROWN [52] enhanced with our PT-LiRPA. As a representative case, we select the MNIST_2x[1024]_ReLU network, which exhibited significant estimation errors in the previous evaluation. The results, reported in Tab 2, show that despite the estimation error, PT-LiRPA produces tighter output bounds than $\alpha$-CROWN, while soundly overapproximating the exact MIP bounds. Similar trends were consistently observed across all evaluated networks, which we do not report here for readability reasons.
@@
For the second question, we consider the models trained on MNIST and CIFAR datasets, as provided in [54]. Hence, we evaluate the performance of our PT-LiRPA-based probabilistic verifier, alongside PROVEN [48], Randomized Smoothing [12], and CROWN [54], by testing for each model 10 random images from the corresponding test datasets. Specifically, we compare the maximum input perturbation $\varepsilon$ that can be certified for each method. CROWN [54] serves as the baseline for worst-case robustness certification, as employed in [48]. Based on the previous results, we set a sample size of $n=350k,\;\xi=0.85$ and $1-2mp\geq 0.99$ in our PT-LiRPA-based verifier. This confidence level is consistent with the settings used by Weng et al. [48] in their experiments.
Certification method Worst-case (CROWN) PROVEN Rand. Smoothing CROWN w/ PT-LiRPA PT-LiRPA certification bound PT-LiRPA certification bound PT-LiRPA certification bound
Confidence  100% $\geq 99\%$ $\geq 99\%$ $\geq 99\%$ increase w.r.t. CROWN increase w.r.t. PROVEN increase w.r.t. Rand. Smoothing
MNIST 2×$\times[1024]$, ReLU 0.03566 0.0556 0.0461 0.0558 1.6X 1.004X 1.21X
MNIST 3×$\times[1024]$, ReLU 0.03112 0.03524 0.03452 0.06652 2.14X 1.9X 1.93X
MNIST 2×$\times[1024]$, Tanh 0.01827 0.02915 0.02301 0.02949 1.61X 1.01X 1.28X
MNIST 3×$\times[1024]$, Tanh 0.01244 0.01360 0.01294 0.0257 2.07X 1.89X 1.99X
MNIST $\times[1024]$, Sigmoid 0.01975 0.02170 0.02439 0.05506 2.79X 2.54X 2.26X
CIFAR 5×$\times[2048]$, ReLU 0.002412 0.00264 0.00778 0.00874 3.62X 3.31X 1.12X
CIFAR 7×$\times[1024]$, ReLU 0.001984 0.00209 0.006264 0.00471 2.37X 2.25X 0X
@@
Table 3: Comparison of PT-LiRPA with worst-case bound CROWN [54] and probabilistic approaches PROVEN [48], Randomized Smoothing [12] on different neural networks MNIST and CIFAR models. $\dagger$ results taken from the original paper [48] due to the code’s unavailability to reproduce the results.
@@
Tab.3 reports the results obtained. The “Worst-case" column indicates the mean of the maximum $\varepsilon$ perturbation tolerated and provably certified using CROWN [54] for that model under consideration in the 10 random images tested, as in [48]. Instead, the PROVEN, Rand. Smoothing and PT-LiRPA columns report the certified $\varepsilon$ we achieve for the same images, using the three probabilistic approaches, sacrificing only a $10^{-2}$ of confidence. In general, we observe that the PT-LiRPA-based verifier can certify robustness levels up to 3.62 times higher than the worst-case baseline CROWN [54] and up to 3.31 and 2.26 times higher compared to PROVEN [48] and Rand. Smoothing [12], respectively. This demonstrates the significant advantage of our approach over other existing probabilistic methods.
Since the worst-case setting employed in this comparison is one of the first LiRPA-based approaches proposed in the literature, in the following we conduct further analysis on the level of tightness we can achieve with respect to more recent LiRPA approaches such as $\alpha$-CROWN [52], $\beta$-CROWN [44], and GCP-CROWN [55]. For each of these approaches, as well as their probabilistic counterparts based on our PT-LiRPA framework, we compute the mean input perturbation $\varepsilon$ that can be certified across 10 random images. As shown in Fig. 6, our probabilistic framework consistently certifies robustness levels at least 1.5 times higher than the worst-case baseline, even when using more recent LiRPA techniques. Notably, the results indicate that as the estimated reachable sets become more precise through over-approximation, the impact of our approach diminishes. Nonetheless, our framework can provide interesting safety information on the model’s robustness level even with very tight provable reachable sets.
@@
Figure 6: Comparison PT-LiRPA with worst-case bound CROWN [54], $\alpha$-CROWN[52], $\beta$-CROWN [44], GCP-CROWN[55] on MNIST 2×$\times[1024]$, ReLU model. On the x-axis, we report the original worst-case method and the corresponding probabilistic version using our PT-LiRPA framework. On the y-axis, we report, for each method, the mean maximum input perturbation $\varepsilon$ that can be certified on 10 random images.
@@
Finally, to assess the computational overhead introduced by the proposed approach, we perform an ablation study analyzing its impact on the total certification time, considering $n=350k$ samples and progressively larger network sizes. We highlight that the overall time complexity of the algorithm is polynomial in the network size, specifically $O\left(n\sum_{i=1}^{N}d_{i}d_{i-1}\right)$, where $d_{i}d_{i-1}$ denotes the number of multiplications at layer $i$. Consequently, the computational time scales linearly with both the number of neurons and the number of samples. This aspect is well highlighted in the results of Tab. 4, where the time overhead of the sampling-based approach in PT-LiRPA is negligible (i.e., less than one second) in the overall certification time, even when employing a significantly large number of samples, thanks to the GPU acceleration employed in the certification process. Clearly, the total computation time of CROWN enhanced with PT-LiRPA is greater than that of using CROWN alone, as the probabilistic certification of a larger tolerable input perturbation results in a longer verification process.
@@
Total certification time Total certification time # computations of interm. bounds # samples Total intem. bounds computation time Mean intem. bounds computation time
MNIST 2×$\times[1024]$, ReLU 35.7s 47.4s 253 350k 0.27s 0.001s
MNIST 3×$\times[1024]$, ReLU 37.73s 54.85s 248 350k 0.34s 0.0014s
MNIST 2×$\times[1024]$, Tanh 23.02s 36.73s 165 350k 0.22s 0.0013s
MNIST 3×$\times[1024]$, Tanh 31.47s 59.82s 186 350k 0.26s 0.0014s
MNIST $\times[1024]$, Sigmoid 40.26s 64.61s 209 350k 0.33s 0.0016s
CIFAR 5×$\times[2048]$, ReLU 26.3s 82.42s 120 350k 0.48s 0.004s
CIFAR 7×$\times[1024]$, ReLU 21.02s 73.9s 154 350k 0.37s 0.0024s
Table 4: Time comparison between CROWN and CROWN enhanced with PT-LiRPA for the certification of the models in Table 3. The Total certification time column reports the overall time required to compute the average $\varepsilon$ perturbation that the model can tolerate across 10 random test images. The # computations of interm. bounds column indicates how many times the intermediate bound computation procedure is invoked to determine the mean $\varepsilon$, while the # samples column specifies the number of samples used in each instance of intermediate bound computation. Finally, the last two columns present the total overhead and the average time per call of the sampling-based approach used to compute the probabilistically optimal intermediate bounds.
@@
To answer the last question, we integrate our PT-LiRPA in the $\alpha,\beta$-CROWN toolbox (https://github.com/Verified-Intelligence/alpha-beta-CROWN) and perform a final experiment on different benchmarks of the VNN-COMP 2022 and 2023 [32, 8]. We point out that the comparison between probabilistic and provable verifiers is performed to have a ground truth (when possible), and to highlight the valuable help that probabilistic approaches can have in solving challenging instances to be verified.
@@
  • ACAS xu [21, 22] benchmark 2023: includes ten properties evaluated across 45 neural networks designed to provide turn advisories for aircraft to prevent collisions. Each neural network consists of 300 neurons distributed over six layers, using ReLU activation functions. The networks take five inputs representing the aircraft’s state and produce five outputs, with the advisory determined by the minimum output value. Here, we verified only property 3, which returns unsafe if the clear of conflict (COC) output is minimal, with a max computation time of 116s.
@@
ACASxu $\alpha$,$\beta$-CROWN $100\%$ 93.33% 42 3 0 26s
$\alpha$,$\beta$-CROWN w/PT-LiRPA $\geq 99\%$ 93.33% 42 3 0 16.37s
tllVerifyBench $\alpha$,$\beta$-CROWN $100\%$ 46.875% 15 17 0 90.2s
$\alpha$,$\beta$-CROWN w/ PT-LiRPA $\geq 99\%$ 46.875% 15 17 0 92s
CIFAR_biasfield $\alpha$,$\beta$-CROWN $100\%$ 95.83% 69 1 2 1553.5s
$\alpha$,$\beta$-CROWN w/ PT-LiRPA $\geq 99\%$ 98.61% 71 1 0 408.7s
CIFAR_tinyimagenet $\alpha$,$\beta$-CROWN $100\%$ 62.5% 15 3 6 1429.6s
$\alpha$,$\beta$-CROWN w/ PT-LiRPA $\geq 99\%$ 87.5% 21 3 0 425.6s
Table 5: Results on VNN-COMP 2022-2023 benchmarks. Results marked in bold report improved performance in terms of verified accuracy (sat instances/all instances) and total verification time for the specific benchmark tested, wr.r.t. a worst-case verification approach.
In Tab. 5 we report our results on the VNN-COMP, where we consider an increased difficulty for the verification process. We start with the simpler benchmark ACASxu [21, 22], and we test property 3. This property is particularly interesting as it holds for 42 of the 45 models tested, thus allowing us to verify the improvement in terms of time and verification accuracy. In the first row of Tab. 5, we can notice that by sacrificing only a $0.01\%$ of confidence, $\alpha,\beta$-CROWN enhanced with PT-LiRPA achieves the same verified accuracy in less verification time, thus confirming our intuition. Interestingly, we observe that tighter bounds are not always beneficial in general. Specifically, in cases where a PGD attack succeeds despite loose bounds, using tighter bounds does not lead to further improvements. Additionally, in some scenarios, less accurate bounds from vanilla LiRPA methods could be quickly refined by BaB, still resulting in efficient verification time. This is exemplified by the tllVerifyBench experiments, where even sacrificing a $0.01\%$ of confidence, PT-LiRPA produced tight estimated reachable sets but achieved the same verified accuracy with a minor overhead in bounds computation.

### v2

@@ 5 Empirical Evaluation
What is the general impact of PT-LiRPA in the verification process of challenging instances such as the one employed in the VNN-COMP (VNN-comp2022; VNN-comp2023)?
All data are collected on a cluster running Rocky Linux 9.34 equipped with Nvidia RTX A6000 (48 GiB) and a CPU AMD Epyc 7313 (16 cores). The code, trained models, and comprehensive instructions for reproducing our results are available at https://github.com/lmarza/ProbVerNet.
@@
To address the first question, we consider the original pre-trained models on MNIST dataset. Specifically, we focus on MLP models with varying depths and activation functions. For consistency and ease of comparison, we adopt the same notation as (crown) and (proven): a model is denoted by the dataset name, followed by the number of layers $i$, the number of neurons per layer $j$, formatted as $i\times[j]$, and the activation function used. We begin by analyzing the mean percentage error in estimating the lower bounds of the intermediate reachable sets (using Eq. 12) for a fixed input image, using PT-LiRPA on various neural networks trained on the MNIST dataset.
@@
Table 1: Mean maximum error in estimating the lower bound of the intermediate reachable set for a fixed input image, using PT-LiRPA with 1p=0.991-p=0.99 on various neural networks trained on the MNIST dataset with different sample sizes.
Name model # samples mean distance # samples mean distance # samples mean distance
error (%) error (%) error (%)
MNIST 2×[$\times[1024]$, ReLU 10000 12.91 100000 9.35 350000 8.78
MNIST 3×[$\times[1024]$, ReLU 10000 9.63 100000 8.49 350000 7.09
MNIST $4]\times[102$4], ReLU 10000 11.89 100000 9.34 350000 8.85
MNIST 2×[$\times[1024]$, Tanh 10000 11.31 100000 12.22 350000 10.49
MNIST 3×[$\times[1024]$, Tanh 10000 10.67 100000 11.01 350000 10.34
MNIST $4]\times[102$4], Sigmoid 10000 5.03 100000 6.21 350000 4.36
@@
The results of Tab. 1 demonstrate that with high confidence (i.e., at least 99%) across all tested networks, increasing the sample size, the percentage error in the estimation decreases. Importantly, even with a limited sample size (e.g., $10000$ samples), we note that the mean error across all the networks in the lower bound estimation is $10.24\%$ from the estimated one, while increasing the sample size, we reach a maximum error of $8.32\%$.
To assess the practical impact of the estimation error on the final output bounds, we consider the same perturbed input of the previous experiments and fix the sample size to $350k$. We then compare the final output bounds obtained by an exact MIP-based verification (MIP), the state-of-the-art $\alpha$-CROWN (acrown) method, and $\alpha$-CROWN (acrown) enhanced with our PT-LiRPA. As a representative case, we select the MNIST_2x[1024]_ReLU network.
[…]
The reason for selecting this network is to enable a comparison with MIP MIP, which provides exact output bounds. Since MIP solvers are inherently designed for ReLU, and more generally, piecewise-linear activations, they are not directly applicable to networks with Sigmoid or Tanh activations. Among the models listed in Table 1, the ones with larger estimation errors under ReLU activations with confidence 1p=0.991-p=0.99 are MNIST_2x[1024]_ReLU and MNIST_4x[1024]_ReLU. We select the former because MIP scalability becomes a limiting factor on larger architectures, making MNIST_2x[1024]_ReLU the most suitable candidate for this analysis. The results, reported in Fig. 7, show that despite the estimation error, PT-LiRPA produces tighter output bounds than $\alpha$-CROWN, while soundly overapproximating the exact MIP bounds. Similar trends, not reported here for the sake of readability, were consistently observed across all evaluated networks compared with $\alpha$-CROWN.
We also perform an additional experiment to analyze the impact of the confidence level on the tightness of the bounds. Specifically, we consider the same model of Fig. 7, i.e., MNIST_2x[1024]_ReLU, where we have the possibility of employing the exact MIP solver and evaluated a range of increasing confidence levels, namely $1-p\in\{0.8,0.9,0.95,0.99,0.995,0.996,0.997,0.998,0.999\}.$ For each confidence level, we compute the bounds using $\alpha$-CROWN, $\alpha$-CROWN enhanced with our PT-LiRPA, and MIP (which provides the exact bounds). To measure the tightness, we calculate for each of the 10 output nodes the distance between the two bounds computed. For example, if MIP returns output bounds [2.2,3.5][-2.2,3.5] and PT-LiRPA returns [2.4,3.7][-2.4,3.7], the distance is given by the sum of the absolute differences between the lower and upper bounds, i.e., |2.2+2.4|+|3.73.5|=0.4|-2.2+2.4|+|3.7-3.5|=0.4. This procedure is repeated for all 10 nodes, collecting, for each confidence level, the mean and standard deviation of these distances.
Figure 8: Mean tightness level of PT-LiRPA for increasing confidence level on MNIST_2x[1024]_ReLU model.
The results in Fig. 8 clearly show that for moderate confidence levels, i.e., $1-p\in[0.8,0.995]$, PT-LiRPA consistently produces sound and tighter bounds compared to $\alpha$-CROWN. As the confidence level approaches 1 (i.e., as $p\to 0$), the distance between PT-LiRPA and the exact bounds increases, leading to looser bounds than $\alpha$-CROWN. This trend is perfectly in line with the probabilistic nature of our approach. In fact, considering, for instance, the formula used to compute the probabilistic lower bound, $\hat{l}=Y_{1}-\frac{Y_{2}-Y_{1}}{(1-p)^{-a_{l}}-1}$ where $a_{l}\approx\frac{\log(\nu)}{\log\left(\frac{Y_{\nu}-Y_{3}}{Y_{3}-Y_{2}}\right)}$ ($\nu=\lfloor n^{\xi}\rfloor=350k^{0.85}\simeq 50k$), we observe that as $p\to 0$ (corresponding to requiring very high confidence), the denominator $(1-p)^{-a_{l}}-1$ tends to 0 faster than the numerator $(Y_{2}-Y_{1})$. As a result, the error term $\frac{Y_{2}-Y_{1}}{(1-p)^{-a_{l}}-1}$ becomes large, and a correspondingly larger margin must be subtracted from the observed minimum. Consequently, the bounds become looser as the confidence level approaches $1$. This behavior is mathematically unavoidable since we cannot guarantee arbitrarily high confidence levels while simultaneously maintaining very tight bounds. The plot therefore not only confirms the soundness of our method but also highlights the expected trade-off between confidence and tightness, which is especially relevant for safety-critical applications. For the sake of completeness, we also tested additional models such as MNIST_2x[1024]_Tanh and MNIST_4x[1024]_Sigmoid. Although MIP was not applicable in these cases, we compared PT-LiRPA only against $\alpha$-CROWN and observed a similar trend, i.e., a loss of tightness above confidence $0.995$, further confirming the generality of our findings.
@@
For the second question, we consider the models trained on MNIST and CIFAR datasets, as provided in (crown). Hence, we evaluate the performance of our PT-LiRPA-based probabilistic verifier, alongside PROVEN (proven), Randomized Smoothing (randomizedSmoothing), and CROWN (crown), by testing for each model 10 random images from the corresponding test datasets. Specifically, we compare the maximum input perturbation $\varepsilon$ that can be certified for each method. CROWN (crown) serves as the baseline for worst-case robustness certification, as employed in (proven). Based on the previous results, we set a sample size of $n=350k,\;\xi=0.85$ and $1-2mp\geq 0.99$ in our PT-LiRPA-based verifier. This confidence level is consistent with the settings used by proven in their experiments.
@@
Table 2: Comparison of PT-LiRPA with worst-case bound CROWN crown and probabilistic approaches PROVEN proven, Randomized Smoothing randomizedSmoothing on different neural networks MNIST and CIFAR models. $\dagger$ results taken from the original paper proven due to the code’s unavailability to reproduce the results.
Certification method Worst-case (CROWN) PROVEN Rand. Smoothing CROWN w/ PT-LiRPA PT-LiRPA certification bound increase
Confidence  100% $\geq 99\%$ $\geq 99\%$ $\geq 99\%$ w.r.t. CROWN, PROVEN, Rand. Smoothing
[…]
MNIST 3×[$\times[1024]$, ReLU 0.03112 $\pm$ 0.01076 $03452±\pm0.0$169 0.06652 $\pm$0.02501 2.14X, 1.9X, 1.93X
[…]
@@
Tab.2 reports the results obtained. The “Worst-case" column indicates the mean and standard deviation of the maximum $\varepsilon$ perturbation tolerated and provably certified using CROWN (crown) for that model under consideration in the 10 random images tested, as in (proven). Instead, the PROVEN, Rand. Smoothing and PT-LiRPA columns report the certified mean $\varepsilon$ and standard deviation we achieve for the same images, using the three probabilistic approaches, sacrificing only a $10^{-2}$ of confidence. In general, we observe that the PT-LiRPA-based verifier can certify robustness levels up to 3.62 times higher than the worst-case baseline CROWN (crown) and up to 3.31 and 2.26 times higher compared to PROVEN (proven) and Rand. Smoothing (randomizedSmoothing), respectively. This demonstrates the significant advantage of our approach over other existing probabilistic methods.
@@
Figure 9: Comparison PT-LiRPA with worst-case bound CROWN (crown), $\alpha$-CROWN(acrown), $\beta$-CROWN (bcrown), GCP-CROWN(GCP-CROWN) on MNIST_2×[$\times[1024]$_ReLU, MNIST_2×[$\times[1024]$_Tanh, and MNIST_4×[$\times[1024]$_Sigmoid models. On the x-axis, we report the original worst-case method and the corresponding probabilistic version using our PT-LiRPA framework. On the y-axis, we report, for each method, the mean maximum input perturbation $\varepsilon$ that can be certified on 10 random images.
Since the worst-case setting employed in this comparison is one of the first LiRPA-based approaches proposed in the literature, we conduct further analysis on the level of tightness we can achieve with respect to more recent LiRPA approaches such as $\alpha$-CROWN (acrown), $\beta$-CROWN (bcrown), and GCP-CROWN (GCP-CROWN). For each of these approaches, as well as their probabilistic counterparts based on our PT-LiRPA framework, we compute the mean input perturbation $\varepsilon$ that can be certified across 10 random images. As shown in Fig. 9, our probabilistic framework consistently certifies robustness levels at least 1.4 times higher than the worst-case baseline, even when using more recent LiRPA techniques. Notably, the results indicate that as the estimated reachable sets become more precise through over-approximation, the impact of our approach diminishes. Nonetheless, our framework can provide interesting safety information on the model’s robustness level even with very tight provable reachable sets.
@@
Finally, to assess the computational overhead introduced by the proposed approach, we perform an ablation study analyzing its impact on the total certification time, considering $n=350k$ samples and progressively larger network sizes. We highlight that the overall time complexity of the algorithm is polynomial in the network size, specifically $O\left(n\sum_{i=1}^{N}d_{i}d_{i-1}\right)$, where the term $\sum_{i=1}^{N}d_{i}d_{i-1}$ represents the complexity for a single sample, where $N$ is the total number of layers in the network, $d_{i-1}$ is the number of neurons in the preceding layer, and $d_{i}$ is the number of neurons in the current layer. The product $d_{i}d_{i-1}$ therefore indicates the number of multiplications required for the processing of layer $i$. Consequently, the total complexity is proportional to the number of samples multiplied by the cost of a single forward propagation through the entire network. This aspect is well highlighted in the results of Tab. 3, where the time overhead of the sampling-based approach in PT-LiRPA is negligible (i.e., less than one second) in the overall certification time, even when employing a significantly large number of samples, thanks to the GPU acceleration employed in the certification process. Clearly, the total computation time of CROWN enhanced with PT-LiRPA is greater than that of using CROWN alone, as the probabilistic certification of a larger tolerable input perturbation results in a longer verification process.
Table 3: Time comparison between CROWN and CROWN enhanced with PT-LiRPA for the certification of the models in Table 2. The Total Cert. Time column reports the overall time required to compute the average $\varepsilon$ perturbation that the model can tolerate across 10 random test images. The # Computations of Interm. Bounds column indicates how many times the intermediate bound computation procedure is invoked to determine the mean $\varepsilon$, while the # Samples column specifies the number of samples used in each instance of intermediate bound computation. Finally, the last two columns present the total overhead and the average time per call of the sampling-based approach used to compute the probabilistically optimal intermediate bounds.
@@
Total Total # Computations # Samples Total interm. bounds Mean Interm. Bounds
Cert. Time Cert. Time of Interm. Bounds Comp. time Comp. time
MNIST 2×[$\times[1024]$, ReLU 35.7s 47.4s 253 350k 0.27s 0.001s
MNIST 3×[$\times[1024]$, ReLU 37.73s 54.85s 248 350k 0.34s 0.0014s
MNIST 2×[$\times[1024]$, Tanh 23.02s 36.73s 165 350k 0.22s 0.0013s
MNIST 3×[$\times[1024]$, Tanh 31.47s 59.82s 186 350k 0.26s 0.0014s
MNIST $4]\times[102$4], Sigmoid 40.26s 64.61s 209 350k 0.33s 0.0016s
CIFAR 5×[$\times[2048]$, ReLU 26.3s 82.42s 120 350k 0.48s 0.004s
CIFAR 7×[$\times[1024]$, ReLU 21.02s 73.9s 154 350k 0.37s 0.0024s
Importantly, the time comparison is conducted only against CROWN as the goal of this experiment is to show that the additional cost introduced by PT-LiRPA is negligible in the overall verification time, while still producing tighter output bounds. Importantly, once the LiRPA baseline is fixed (e.g., $\alpha$-CROWN, $\beta$-CROWN, GCP-CROWN), the subsequent linearization procedure is identical whether using the original method or our PT-LiRPA variant. The only difference lies in the way intermediate bounds are computed, which are then used to construct the diagonal matrices and bias terms. Consequently, we argue that measuring the overhead against CROWN is representative, since the additional cost introduced by PT-LiRPA does not depend on which LiRPA baseline is employed.
@@
To answer the last question, we integrate our PT-LiRPA in the $\alpha,\beta$-CROWN toolbox (https://github.com/Verified-Intelligence/alpha-beta-CROWN) and perform a final experiment on different benchmarks of the VNN-COMP 2022 and 2023 (VNN-comp2022; VNN-comp2023). We point out that the comparison between probabilistic and provable verifiers is performed to have a ground truth (when possible), and to highlight the valuable help that probabilistic approaches can have in solving challenging instances to be verified.
@@
  • ACAS xu (ACAS; Reluplex) benchmark 2023: includes ten properties evaluated across 45 neural networks designed to provide turn advisories for aircraft to prevent collisions. Each neural network consists of 300 neurons distributed over six layers, using ReLU activation functions. The networks take five inputs representing the aircraft’s state and produce five outputs, with the advisory determined by the minimum output value. Here, we verified only property 3, which returns unsafe if the clear of conflict (COC) output is minimal, with a max computation time of 116s.
@@
Table 4: Results on VNN-COMP 2022-2023 benchmarks. Results marked in bold report improved performance in terms of verified accuracy (% sat instances/all instances) and total verification time for the specific benchmark tested, w.r.t. a worst-case verification approach.
@@
ACASxu $\alpha$$\beta$-CROWN $100\%$ 93.33% 42 3 0 26s
$\alpha$$\beta$-CROWN
w/PT-LiRPA $\geq 99\%$ 93.33% 42 3 0 16.37s
tllVerifyBench $\alpha$$\beta$-CROWN $100\%$ 46.875% 15 17 0 90.2s
$\alpha$$\beta$-CROWN
w/PT-LiRPA $\geq 99\%$ 46.875% 15 17 0 92s
CIFAR_biasfield $\alpha$$\beta$-CROWN $100\%$ 95.83% 69 1 2 1553.5s
$\alpha$$\beta$-CROWN
w/PT-LiRPA $\geq 99\%$ 98.61% 71 1 0 408.7s
CIFAR_tinyimagenet $\alpha$$\beta$-CROWN $100\%$ 62.5% 15 3 6 1429.6s
$\alpha$$\beta$-CROWN
w/PT-LiRPA $\geq 99\%$ 87.5% 21 3 0 425.6s
In Tab. 4 we report our results on the VNN-COMP, where we consider an increased difficulty for the verification process. We start with the simpler benchmark ACASxu (ACAS; Reluplex), and we test property 3. This property is particularly interesting as it holds for 42 of the 45 models tested, thus allowing us to verify the improvement in terms of time and verification accuracy. In the first row of Tab. 4, we can notice that by sacrificing only a $0.01\%$ of confidence, $\alpha,\beta$-CROWN enhanced with PT-LiRPA achieves the same verified accuracy in less verification time, thus confirming our intuition. Interestingly, we observe that tighter bounds are not always beneficial in general. Specifically, in cases where a PGD attack succeeds despite loose bounds, using tighter bounds does not lead to further improvements. Additionally, in some scenarios, less accurate bounds from vanilla LiRPA methods could be quickly refined by BaB, still resulting in efficient verification time. This is exemplified by the tllVerifyBench experiments, where even sacrificing a $0.01\%$ of confidence, PT-LiRPA produced tight estimated reachable sets but achieved the same verified accuracy with a minor overhead in bounds computation.
@@
6 Assumptions and Limitations
The proposed PT-LiRPA framework builds on several assumptions that define its applicability and theoretical guarantees. Our probabilistic framework relies on uniform random sampling within the perturbation region $\mathcal{C}_{\mathbf{x}_{0},\epsilon}$ to estimate probabilistically tight reachable sets. Consequently, our theoretical probabilistic guarantees, derived from wilks’s tolerance limit theorem and extended using extreme value theory (haan2006extreme), hold under the assumption that the samples are independent and representative of the true input distribution within $\mathcal{C}$.
Importantly, by accepting robustness certificates that hold for a fraction $R$ of the perturbation region, the theoretical and practical tool derived from wilks’s results, i.e., Theorem 3.4, remains a valuable solution, as it provides a closed-form expression to compute the number of samples required for any desired confidence level $\psi$ and coverage ratio $R$. To address this limitation and extend the analysis to probabilistic certificates valid over the entire perturbation region, thus aligning with the probabilistic verification literature, we further based our theoretical derivation on extreme value theory and the results of de1981estimation. In this case, our derivation in Theorem 3.7 provides, for any choice of sample size $n$ and confidence error $p$, a guarantee that holds for all input values ${\bm{x}}\in\mathcal{C}$. However, unlike Wilks’ approach, EVT does not yield a closed-form expression to determine the minimum number of samples required to achieve a desired precision and confidence level. Consequently, the balance between probabilistic soundness and the tightness of the resulting output bounds becomes more empirical, as confirmed in our experiments. In fact, the tightness of the computed bounds and the computational efficiency of the method depend on the chosen sample size $n$, and the estimate of the tail distribution $\nu=\lfloor n^{\xi}\rfloor$, which must balance verification accuracy and computational cost. Despite these limitations, we argue that PT-LiRPA complements deterministic verification methods by offering practical, quantifiable robustness guarantees for challenging instances where worst-case formal verification is either overly conservative or computationally infeasible.